OpenStudy (anonymous):

please help solve ( x - 5)(x + 10) =0 this is a quadratic equation do not know how to work the problem

OpenStudy (amistre64):

if a*b = 0, then what is true of a or b?

OpenStudy (anonymous):

are you saying 0x5 =0?

OpenStudy (amistre64):

in a sense yes. but lets be more precise: a*b = 0 is true when a=0: 0*b = 0 and when b=0: a*0 = 0 and when both are 0: 0*0 = 0 remember how pointless if seemed to do the zero times tables?

OpenStudy (amistre64):

( x - 5)(x + 10) =0 is true when: x - 5 = 0 or x + 10 = 0 solve each instance for x

OpenStudy (anonymous):

Yes I do...I have been out of school 20 something years and this algebra is killing me I just cannot get it.

OpenStudy (amistre64):

it was about 20 years for me too before i tried college

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

so knowing that either part can be zero, how would we solve the following for x x - 5 = 0 or x + 10 = 0

OpenStudy (anonymous):

x would be 1 or 5?

OpenStudy (anonymous):

then again it could be 0

OpenStudy (amistre64):

if its not obvious, we can work the algebra .... x - 5 = 0 ; add 5 to both sides to get x all by itself +5 +5 --------- x = 5 x + 10 = 0 ; subtract 10 from both sides to get x all by itself -10 -10 ---------- x = -10

OpenStudy (amistre64):

lets test these out ( x - 5)(x + 10) = 0 , let x=5 ( 5 - 5)(5 + 10) = 0 0(15) = 0 ; true --------------------------- ( x - 5)(x + 10) = 0 , let x=-10 (-10- 5)(-10+ 10) = 0 -15(0) = 0 ; true

OpenStudy (anonymous):

so that is the way I need to work the problem correct?